OpenStudy (anonymous):

pls help The length of a field is three times its width. If the perimeter is 64 m, what is the width of the field? A. 8 m B. 16 m C. 21 m D. 32 m

OpenStudy (anonymous):

i think it's A

OpenStudy (alexandervonhumboldt2):

l=3w w 2*(l+w)=64 l+w=32 3w+w=32 4w=32 w=8 thus you are correct

OpenStudy (anonymous):

Ty. :D

OpenStudy (youngstudier):

Good job, Starz :)

OpenStudy (anonymous):

Thank you :D :3